OpenStudy (anonymous):

(8x)/(x^2+4x+4)/(4x^2)/(x^2-4) in simplified form???

OpenStudy (anonymous):

\[\frac{8x}{x^2+4x+4}\times \frac{x^2-4}{4x^2}\]

OpenStudy (anonymous):

we can cancel right away a bit, but can also factor

OpenStudy (anonymous):

\[\frac{8x(x-2)(x+2)}{4x^2(x+2)(x+2)}\]

OpenStudy (anonymous):

cancel away and get \[\frac{2(x-2)}{x(x+2)}\]

OpenStudy (anonymous):

Wow thank you so much :) Can I ask you about one more? x^2-36 /x+6 I think it's x-6 but I don't know what the restriction would be?

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

What would the restriction be?

OpenStudy (anonymous):

6 or -6?

OpenStudy (anonymous):

x cannot be -6